# Chunk Based Speech Pre-training with High Resolution Finite Scalar Quantization

### v1

@@ 5.3 Latency evaluation
We use the base Transducer model to analyze the impact of different chunk sizes on latency and performance. We use Length-Adaptive Average Lagging(LAAL) (Papi et al., 2022) as latency scorer, which is evaluated with the Simuleval toolkit (Ma et al., 2020). We employ greedy decoding instead of beam search decoding for simplicity in this study.
Figure 3 depicts the model performance (blue) and latency (red) under different chunk sizes for speech recognition (3(a)) and speech translation (3(b) and 3(c)). The speech recognition is evaluated on the Librispeech dev-other set and speech translation tasks are evaluated on the corresponding MuST-C dev set. It can be seen that the model is with a small latency (LAAL 0.47 second) but high WER (7.6) when the chunk size is small (160ms). The performance improves steadily when the chunk size increases but at the cost of increasing latency. The speech translation experiments have similar observations. The results indicate the model can switch from streaming mode to offline mode smoothly as the chunk sizes change. Other latency evaluation results, such as Average Lagging (AL) (Ma et al., 2019) and Differentiable Average Lagging (DAL) (Arivazhagan et al., 2019), could be found in Appendix B
@@ 6 Related Work
Speech self-supervised learning algorithms form contextual representations through learning to predict missing information due to time constraint or masking. van den Oord et al. (2018) propose an auto-regressive based SSL approach to predict the future samples based on the previous samples. Speech SSL methods such as wav2vec2 (Baevski et al., 2020b), HuBert (Hsu et al., 2021b) and BEST-RQ (Chiu et al., 2022) choose to randomly mask audio span and the model learns to restore those masked regions. Chunk SSL combines both approaches that it masks audio span in the extended chunk and recovers masking frame indices in an chunk based auto-regressive manner.
Vector quantization (VQ) is one of the important components in the speech SSL. Baevski et al. (2020a) leverages Gumbel-Softmax with multiple variable groups to discretize audio features. Methods such as diversity loss have to take to alleviate codeword collapse issue. Hsu et al. (2021b) resorts k-means to cluster the encoder outputs to build the codebook. In order to achieve a good representation, an iterative training is required to keep re-building the encoder with a new codebook. BEST-RQ (Chiu et al., 2022) chooses an alternative way, which initializes the codebook randomly and no further update is applied during SSL. Compared with these VQ methods aforementioned, FSQ does not suffer from codebook collapse and achieves high codebook usage (Mentzer et al., 2024).
@@ A.2 Chunk SSL pre-training setting
The pre-training is scheduled as Transformer (Vaswani et al., 2017) with warmup step 25,000 and maximum learning rate 3e-4. The batch size is 16,000 seconds for the base configure model. We choose a smaller effective batch size for the large configure model with 6,000 seconds due to computation resource limitation. The model is trained up to 400,000 updates. It takes about 10 days for 8 A100 GPUs to build a Chunk SSL encoder with the base configure.
@@ A.3 Librispeech fine-tuning setting
960 hours fine-tuning For the base configure model, we choose the three stages scheduler scheme as (Baevski et al., 2020b), i.e., warmup, hold and annealing. The warmup stage takes about 1000 updates with encoder parameters frozen, the hold stage takes about 40,000 updates and the remaining 60,000 updates are in an annealing stage. The peak learning rate in fine-tuning is 2.0e-3. The SpecAugment (Park et al., 2019) data augmentation is applied with 2 frequency maskings of 27, and 10 time maskings at a maximum ratio of 0.05 in fine-tuning experiments. The effective batch size for fine-tuning is 10,000 seconds. For the large configure model, we choose Transformer scheduler (Vaswani et al., 2017) since it has less hype-parameters to explore. The warmup step is 5000 with encoder parameters frozen. Similar as the base configure model, the total update step is up to 100,000. The peak learning rate is 5e-4. The effective batch size for the large model is 7200 seconds.
@@ A.4 MuST-C fine-tuning settings
The translation Transducer model is optimized with Transformer (Vaswani et al., 2017) scheduler with warmup step 10,000 and total update step up to 100,000. The peak learning rate is 5e-4. The effective batch size is 6400 seconds.
@@ Appendix B Latency Results
The full latency evaluation results including Average Lagging (AL) (Ma et al., 2019), Length Adaptive Average Lagging (LAAL) (Papi et al., 2022), Average Proportion (AP) (Cho & Esipova, 2016), Differentiable Average Lagging (DAL) (Arivazhagan et al., 2019), and Average Token Delay (ATD) (Kano et al., 2023) for the Librispeech and Must-C are presented in Table 5, Table 6 and Table 7.

### v2

@@ 5.4 Latency evaluation
We use the base Transducer model to analyze the impact of different chunk sizes on latency and performance. We use Length-Adaptive Average Lagging (LAAL) (Papi2022OverGenerationCB) as latency scorer, which is evaluated with the Simuleval toolkit (ma2020simuleval). We employ greedy decoding instead of beam search decoding for simplicity in this study.
Figure 3 depicts the model performance (blue) and latency (red) under different chunk sizes for speech recognition (3(a)) and speech translation (3(b) and 3(c)). The speech recognition is evaluated on the Librispeech dev-other set and speech translation tasks are evaluated on the corresponding MuST-C dev set. It can be seen that the model is with a small latency (LAAL 0.47 second) but high WER (7.6) when the chunk size is small (160ms). The performance improves steadily when the chunk size increases but at the cost of increasing latency. The speech translation experiments have similar observations. The results indicate the model can switch from streaming mode to offline mode smoothly as the chunk sizes change. Other latency evaluation results, such as Average Lagging (AL) (Ma2018STACLST) and Differentiable Average Lagging (DAL) (Arivazhagan2019MonotonicIL), could be found in Appendix B
@@ 6 Related Work
Speech self-supervised learning algorithms form contextual representations through learning to predict missing information due to time constraint or masking. Oord2018RepresentationLW propose an auto-regressive based SSL approach to predict the future samples based on the previous samples. Speech SSL methods such as wav2vec2 (Baevski2020wav2vecV2), HuBert (Hsu2021HubertHM) and BEST-RQ (Chiu2022SelfsupervisedLW) choose to randomly mask audio span and the model learns to restore those masked regions. Chunk SSL combines both approaches that it masks audio span in the extended chunk and recovers masking frame indices in an chunk based auto-regressive manner.
Vector quantization (VQ) is one of the important components in the speech SSL. Baevski2020vqwav2vecSL leverages Gumbel-Softmax with multiple variable groups to discretize audio features. Methods such as diversity loss have to take to alleviate codeword collapse issue. Hsu2021HubertHM resorts k-means to cluster the encoder outputs to build the codebook. In order to achieve a good representation, an iterative training is required to keep re-building the encoder with a new codebook. BEST-RQ (Chiu2022SelfsupervisedLW) chooses an alternative way, which initializes the codebook randomly and no further update is applied during SSL. Compared with these VQ methods aforementioned, FSQ does not suffer from codebook collapse and achieves high codebook usage (Mentzer2023FiniteSQ).
@@ A.2 Chunk SSL pre-training setting
The pre-training is scheduled as Transformer (Vaswani2017AttentionIA) with warmup step 25,000 and maximum learning rate 3e-4. The batch size is 16,000 seconds for the base configuration model. We choose a smaller effective batch size for the large configuration model with 6,000 seconds due to computational resource limitation. The model is trained up to 400,000 updates. It takes about 10 days for 8 A100 GPUs to build a Chunk SSL encoder with the base configuration.
@@ A.3 Librispeech fine-tuning setting
960 hours of fine-tuning For the base configuration model, we choose the three stages scheduler scheme as (Baevski2020wav2vecV2), i.e., warmup, hold and annealing. The warmup stage takes about 1000 updates with encoder parameters frozen, the hold stage takes about 40,000 updates and the remaining 60,000 updates are in an annealing stage. The peak learning rate in fine-tuning is 2.0e-3. The SpecAugment (park2019specaugment) data augmentation is applied with 2 frequency maskings of 27, and 10 time maskings at a maximum ratio of 0.05 in fine-tuning experiments. The effective batch size for fine-tuning is 10,000 seconds. For the large configure model, we choose Transformer scheduler (Vaswani2017AttentionIA) since it has less hype-parameters to explore. The warmup step is 5000 with encoder parameters frozen. Similar as the base configure model, the total update step is up to 100,000. The peak learning rate is 5e-4. The effective batch size for the large model is 7200 seconds.
10 hours fine-tuning We present the low resource results with 10 hours fine-tuning data at Appendix C. The model is optimized with a two-stage scheduler with warmup and hold stages. The warmup and hold steps are 2000 and 8000 for both base and large configures. The encoder is frozen during the warmup stage. Small peak learning rate 8e-05 and 3e-5 are used for the large and base models respectively. We use different SpecAugment parameters for the 10 hours fine-tuning with 1 frequency maskings of 34, and 10 time maskings at a maximum ratio of 0.02.
@@ A.4 MuST-C fine-tuning settings
The translation Transducer model is optimized with Transformer (Vaswani2017AttentionIA) scheduler with warmup step 10,000 and total update step up to 100,000. The peak learning rate is 5e-4. The effective batch size is 6400 seconds.
@@ Appendix B Latency Results
The full latency evaluation results including Average Lagging (AL) (Ma2018STACLST), Length Adaptive Average Lagging (LAAL) (Papi2022OverGenerationCB), Average Proportion (AP) (Cho2016CanNM), Differentiable Average Lagging (DAL) (Arivazhagan2019MonotonicIL), and Average Token Delay (ATD) (Kano2022AverageTD) for the Librispeech and Must-C are presented in Table 6, Table 7 and Table 8.
Appendix C Low resource
| Model | dev-clean | dev-other | test-clean | test-other | ave. |
| wav2vec2 Base (Baevski2020wav2vecV2) | 10.9 | 17.4 | 11.1 | 17.6 | 14.3 |
| wav2vec2 Large (Baevski2020wav2vecV2) | 6.3 | 9.8 | 6.3 | 10.0 | 8.1 |
| Chunk SSL Base (O) | 10.9 | 18.6 | 10.9 | 19.2 | 14.9 |
| Chunk SSL Base (S) | 12.0 | 21.8 | 12.0 | 22.3 | 17.0 |
| Chunk SSL Large (O) | 9.6 | 15.3 | 10.0 | 15.9 | 12.7 |
| Chunk SSL Large (S) | 10.5 | 18.5 | 10.9 | 18.9 | 14.7 |
The self-supervised trained model can leverage large amounts of unlabeled data to build an universal representation, which makes it possible to build a speech recognition model with relatively small amounts of labeled data. We examine Chunk SSL pre-trained models with 10 hours Librispeech data (Kahn2019LibriLightAB). We chose a small predictor, which is with input embedding size 128 and forward layer dimension 512, due to limited amount of data. The results are demonstrated in Table 9, where “O” and “S” stand for offline and streaming recognition respectively. According to Table 9, Chunk SSL models achieve compelling results even with just 10 hours of data. The base configure model obtains similar or slightly better offline results compared with the wav2vec2 base model in the two clean datasets, and about 1 WER worse for the two other datasets. The Chunk SSL large model reduces the WER further compared with the base configure model, though they are still significant behind the corresponding wav2vec2 large model. We believe it is due to limited computation resource for the large model pre-training. We have to choose a smaller batch size for the large model and re-use the same configure from the base model pre-training. On the other hand, the streaming results are promising that the streaming model only increase the WER by less than 20% compared with the corresponding offline results. It shows the effectiveness of the proposed Chunk SSL pre-training.